A footballer kicks a ball with a force of 500N if the ball meets his leg just for 1 sec the impulse acted on the ball is
1
Expert's answer
2020-04-23T12:45:39-0400
"I=F\Delta t"

"I=(500)(1)=500\frac{kg\cdot m}{s}"